### v1

@@ 4.1 Theoretical results for random latent positions
The model in Section 2 and the analysis in Section 4 are for deterministic latent positions $\{\mathbf{X}^{(i)}_{t}\}$. Actually, the theoretical guarantees can also be provided for a case where $\{\mathbf{X}^{(i)}_{t}\}$ are randomly generated from latent position stochastic processes as described in [8].
@@
Definition 4.7 (Inner product distribution).
@@
For multiple dynamic networks $\{G^{(i)}_{t}\}_{i\in[m],t\in[T]}$, we denote the corresponding latent position stochastic processes by $\{X^{(i)}\}$. For each fixed $i\in[m]$, as in the case of deterministic latent positions, we define the distance between any two latent position matrices $\mathbf{X}^{(i)}_{t_{1}}$ and $\mathbf{X}^{(i)}_{t_{2}}$ in Eq. (1) for different time points, here we also define a distance to quantify the difference between two latent position distributions $X_{t_{1}}$ and $X_{t_{2}}$. For any pair of time points $t_{1}$ and $t_{2}$ in $[T]$, we define a distance between the two latent position distributions $X^{(i)}_{t_{1}}$ and $X^{(i)}_{t_{2}}$ as
(6) $\breve{\mathbf{D}}^{(i)}_{t_{1},t_{2}}:=\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbb{E}[(\mathbf{O}X^{(i)}_{t_{1}}-X^{(i)}_{t_{2}})(\mathbf{O}X^{(i)}_{t_{1}}-X^{(i)}_{t_{2}})^{\top}]\|^{1/2},$
@@
For the multiple dynamic networks clustering problem of interest, the $m$ dynamic networks are divided into $K$ clusters with cluster labels $\{Y_{i}\}_{i\in[m]}$. We suppose the dynamic networks in the same cluster generated from the same latent position stochastic process. That is, for any $i,j\in[m]$ such that $Y_{i}=Y_{j}$, $X^{(i)}$ and $X^{(j)}$ follow the same distribution. Then with Eq. (6) we have $\breve{\mathbf{D}}^{(i)}=\breve{\mathbf{D}}^{(j)}$, and further according to Theorem 2.6, if $\lambda_{r}(\breve{\mathbf{B}}^{(i)})>\lambda_{r+1}(\breve{\mathbf{B}}^{(i)})$, the mirrors $\breve{\mathbf{M}}^{(i)}$ and $\breve{\mathbf{M}}^{(j)}$ are the same up to an $r\times r$ orthogonal matrix. Recall that we are more interested in the situation where mirrors of different clusters can be distinguished. We suppose there exists a distinct parameter $\breve{c}>0$ such that for any two dynamic networks $i,j\in[m]$ with $Y_{i}\neq Y_{j}$, we have $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\breve{\mathbf{M}}^{(i)}\mathbf{O}-\breve{\mathbf{M}}^{(j)}\|_{F}\geq\breve{c}.$ We define the distance between mirrors as $\breve{\mathbf{D}}^{\star}_{i,j}=\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\breve{\mathbf{M}}^{(i)}\mathbf{O}-\breve{\mathbf{M}}^{(j)}\|_{F}$, and then have the pairwise mirror distance matrix $\breve{\mathbf{D}}^{\star}$ with the clear block structure as shown in Eq. (3).
We now analyze the theoretical results of DNCMD under this model with randomly generated latent positions. The error bounds for $\hat{\mathbf{M}}^{(i)}$ and $\hat{\mathbf{D}}^{\star}$ as the estimates of $\breve{\mathbf{M}}^{(i)}$ and $\breve{\mathbf{D}}^{\star}$ is given in Theorem 4.8 and Theorem 4.9, respectively. We want to emphasize that, under the assumption that for each dynamic network, the latent positions of all vertices are i.i.d. samples of a stochastic process, it is natural to allow different dynamic networks to have different numbers of vertices. Specifically, we can suppose dynamic network $i$ has $n_{i}$ vertices for all $i\in[m]$, and Algorithm 1 can be extended to such a scenario. For the following theoretical results, we let $n=\min_{i\in[m]}\{n_{i}\}$.
Theorem 4.8.
@@
Theorem 4.9.
Consider the setting of Theorem 4.8. We suppose there exists a parameter $\breve{\lambda}$ such that $\lambda_{r}(\breve{\mathbf{B}}^{(i)})\asymp\breve{\lambda}$ for all $i\in[m]$. We then have $\|\hat{\mathbf{D}}^{\star}-\breve{\mathbf{D}}^{\star}\|_{\max}\lesssim\frac{T^{1/2}(n\breve{\rho}_{n})^{1/2}\log n}{n\breve{\lambda}^{1/2}}$ with high probability. We further suppose the distinct parameter $\breve{c}$ satisfies
@@
Then Eq. (5) for $\hat{\mathbf{D}}^{\star}$ holds, and hierarchical clustering for $\hat{\mathbf{D}}^{\star}$ achieves exact recovery.
Recall that we suppose $\|\mathbb{E}[X^{(i)}_{t}X^{(i)\top}_{t}]\|\asymp\breve{\rho}_{n}$ with some sparsity parameter $\breve{\rho}$. If we further suppose there exists a factor $\breve{\tilde{\rho}}>0$ to describe the magnitude of the change of $\{X^{(i)}_{t}\}_{t\in[T]}$ such that $\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbb{E}[(\mathbf{O}X^{(i)}_{t_{1}}-X^{(i)}_{t_{2}})(\mathbf{O}X^{(i)}_{t_{1}}-X^{(i)}_{t_{2}})^{\top}]\|\asymp\breve{\rho}_{n}\breve{\tilde{\rho}}$ for $t_{1}\neq t_{2}$, the similar analysis as Proposition 4.4 and Proposition 4.5 can be derived.
@@ 5.1 Error in mirror estimate
We demonstrate the theoretical results for the estimated mirror in Theorem 4.2 with numerical simulation. Theorem 4.2 states a result for any fixed dynamic network $i\in[m]$. For ease of exposition, in this section we will fix a value of $i\in[m]$ and thereby drop the index $i$ from our matrices. We generate the true latent positions $\{\mathbf{X}_{t}\}$ with the following random walk process. Recall that the true latent position of vertex $s\in[n]$ in time point $t\in[T]$ is the $s$th row of $\mathbf{X}_{t}$. Let $d=1$ and denote the track of the latent position of vertex $s\in[n]$ as $\{X^{<s>}_{t}\in\mathbb{R}\}_{t\in[T]}$. Let $\tilde{c}\geq 0,\delta_{T}>0$ be two constants satisfying $\tilde{c}+\delta_{T}T\leq 1$. For a fixed $p\in(0,1)$, for all $s\in[n]$, we generate $\{X^{<s>}_{t}\in\mathbb{R}\}_{t\in[T]}$ as follows independently.
(7) $X^{<s>}_{0}=\tilde{c},\text{ and for }t\in[T],X^{<s>}_{t}=\begin{cases}X^{<s>}_{t-1}+\delta_{T}&\text{ with probability }p,\\ X^{<s>}_{t-1}&\text{ with probability }1-p.\end{cases}$
When we set $\delta_{T}=(1-\tilde{c})/T$, an observation is that, the dynamic network with the latent positions generated from this model has $\mathbf{B}$ with $\lambda_{1}(\mathbf{B})\asymp T$ and $\lambda_{k}(\mathbf{B})\lesssim 1$ for $k=2,\dots,T$; see [13] for more details, and then we know $r=1$. We demonstrate the error rate of $\hat{\mathbf{M}}$ in Theorem 4.2 under this setting. More specifically, we fix $\tilde{c}=0.1,p=0.4$, and either fix $n=100$ and vary $T\in\{20,30,40,50,70,100\}$ or fix $T=10$ and vary $n\in\{50,100,200,400,800\}$. The estimation error of $\hat{\mathbf{M}}$ is evaluated using $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}\mathbf{O}-\mathbf{M}\|_{F}$, and the results are summarized in Fig. 2. For this setting, according to Theorem 4.2 and Proposition 4.4, if we only vary $T$, we have $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}\mathbf{O}-\mathbf{M}\|_{F}\lesssim 1$, and if we only vary $n$, we have $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}\mathbf{O}-\mathbf{M}\|_{F}\lesssim n^{-1/2}$. We see the changes of $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}\mathbf{O}-\mathbf{M}\|_{F}$ in Fig. 2 comply with the theoretical error rates obtained from Theorem 4.2 and Proposition 4.4.
@@ 5.2 Clustering performance
We perform a numerical experiment to show the clustering performance of DNCMD. We consider $K=2$ clusters of the dynamic networks generated with Eq. (7). These two clusters each have $20$ dynamic networks, and thus we need to cluster $m=2\times 20=40$ dynamic networks. For the dynamic networks, we fix the number of time points $T=50$, and the middle time point $t=25$ is a change point of the probability of random walk $p$. More specifically, for cluster $1$, we set $p=0.45$ for $t=1,\dots,25$ and set $p=0.55$ for $t=26,\dots,50$. And for cluster $2$, we set $p=0.55$ for $t=1,\dots,25$ and set $p=0.45$ for $t=26,\dots,50$. According to Theorem 4.3, Proposition 4.5, and Theorem 4.9, we know DNCMD achieves exact recovery, and it means that for $n$ large enough DNCMD can recover the true clusters with high probability. We vary $n\in\{20,30,40,80,120,200\}$ to see the clustering performance of DNCMD.
To measure the clustering accuracy, we compute the adjusted rank index (ARI) for the similarity between the estimated cluster labels and the true labels. The higher the ARI value, the closer the empirical clustering result and the true label assignment are to each other. ARI ranges from $-1$ to $1$, where $1$ indicates perfect agreement between the empirical clustering result and the true label assignment, and 00 indicates a random agreement. Since there is no existing clustering algorithm specifically designed for such dynamic network clustering problems, we compare the results of DNCMD with those of the classical clustering algorithm, k-means. More specifically, we consider using k-means to directly cluster the original adjacency matrices $\{\mathbf{A}^{(i)}\}$, or using k-means to cluster the estimated low-rank latent position matrices $\{\hat{\mathbf{X}}^{(i)}\}$. We run $100$ independent Monte Carlo replicates and summarize the result in Fig. 3. Fig. 3 shows that, when $n$ is large enough (in this example, $n$ only needs to exceed about $100$), DNCMD always has the ARI very close to $1$ and can recover the true clusters perfectly. And the k-means algorithm applied to the original adjacency matrices or the latent position matrices is not very effective and stable for this problem.
@@ 6 Real Data Experiments
In this section, we cluster the trade dynamic networks to analyze the similarities and differences in trade evolution across different food products in Section 6.1, and demonstrate the use of DNCMD on Drosophila larval connectome data to analyze edge functions in Section 6.2.

### v2

@@ 4.1 Theoretical results for random latent positions
The model in Section 2 and the analysis in Section 4 are for deterministic latent positions $\{\mathbf{X}^{(i)}_{t}\}$. Actually, the theoretical guarantees can also be provided for a case where $\{\mathbf{X}^{(i)}_{t}\}$ are randomly generated from latent position stochastic processes as described in Athreya et al. (2024).
@@
Definition 4 (Inner product distribution).
@@
For multiple dynamic networks $\{G^{(i)}_{t}\}_{i\in[m],t\in[T]}$, we denote the corresponding latent position stochastic processes by $\{X^{(i)}\}$. For each fixed $i\in[m]$, as in the case of deterministic latent positions, we define the distance between any two latent position matrices $\mathbf{X}^{(i)}_{t_{1}}$ and $\mathbf{X}^{(i)}_{t_{2}}$ in Eq. (2.1) for different time points, here we also define a distance to quantify the difference between two latent position distributions $X_{t_{1}}$ and $X_{t_{2}}$. For any pair of time points $t_{1}$ and $t_{2}$ in $[T]$, we define a distance between the two latent position distributions $X^{(i)}_{t_{1}}$ and $X^{(i)}_{t_{2}}$ as
$\breve{\mathbf{D}}^{(i)}_{t_{1},t_{2}}:=\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbb{E}[(\mathbf{O}X^{(i)}_{t_{1}}-X^{(i)}_{t_{2}})(\mathbf{O}X^{(i)}_{t_{1}}-X^{(i)}_{t_{2}})^{\top}]\|^{1/2},$ (4.3)
@@
For the multiple dynamic networks clustering problem of interest, the $m$ dynamic networks are divided into $K$ clusters with cluster labels $\{Y_{i}\}_{i\in[m]}$. We suppose the dynamic networks in the same cluster generated from the same latent position stochastic process. That is, for any $i,j\in[m]$ such that $Y_{i}=Y_{j}$, $X^{(i)}$ and $X^{(j)}$ follow the same distribution. Then with Eq. (4.3) we have $\breve{\mathbf{D}}^{(i)}=\breve{\mathbf{D}}^{(j)}$, and further according to Theorem 2, if $\lambda_{r}(\breve{\mathbf{B}}^{(i)})>\lambda_{r+1}(\breve{\mathbf{B}}^{(i)})$, the mirrors $\breve{\mathbf{M}}^{(i)}$ and $\breve{\mathbf{M}}^{(j)}$ are the same up to an $r\times r$ orthogonal matrix. Recall that we are more interested in the situation where mirrors of different clusters can be distinguished. We suppose there exists a distinct parameter $\breve{c}>0$ such that for any two dynamic networks $i,j\in[m]$ with $Y_{i}\neq Y_{j}$, we have $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\breve{\mathbf{M}}^{(i)}\mathbf{O}-\breve{\mathbf{M}}^{(j)}\|_{F}\geq\breve{c}.$ We define the distance between mirrors as $\breve{\mathbf{D}}^{\star}_{i,j}=\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\breve{\mathbf{M}}^{(i)}\mathbf{O}-\breve{\mathbf{M}}^{(j)}\|_{F}$, and then have the pairwise mirror distance matrix $\breve{\mathbf{D}}^{\star}$ with the clear block structure as shown in Eq. (3.1).
We now analyze the theoretical results of DNCMD under this model with randomly generated latent positions. The error bounds for $\hat{\mathbf{M}}^{(i)}$ and $\hat{\mathbf{D}}^{\star}$ as the estimates of $\breve{\mathbf{M}}^{(i)}$ and $\breve{\mathbf{D}}^{\star}$ is given in Theorem 6 and Theorem 7, respectively. We want to emphasize that, under the assumption that for each dynamic network, the latent positions of all vertices are i.i.d. samples of a stochastic process, it is natural to allow different dynamic networks to have different numbers of vertices. Specifically, we can suppose dynamic network $i$ has $n_{i}$ vertices for all $i\in[m]$, and Algorithm 1 can be extended to such a scenario. For the following theoretical results, we let $n=\min_{i\in[m]}\{n_{i}\}$.
Theorem 6.
@@
Theorem 7.
Consider the setting of Theorem 6. We suppose there exists a parameter $\breve{\lambda}$ such that $\lambda_{r}(\breve{\mathbf{B}}^{(i)})\asymp\breve{\lambda}$ for all $i\in[m]$. We then have $\|\hat{\mathbf{D}}^{\star}-\breve{\mathbf{D}}^{\star}\|_{\max}\lesssim\frac{T^{1/2}(n\breve{\rho}_{n})^{1/2}\log n}{n\breve{\lambda}^{1/2}}$ with high probability. We further suppose the distinct parameter $\breve{c}$ satisfies
@@
Then Eq. (4.2) for $\hat{\mathbf{D}}^{\star}$ holds, and hierarchical clustering for $\hat{\mathbf{D}}^{\star}$ achieves exact recovery.
Recall that we suppose $\|\mathbb{E}[X^{(i)}_{t}X^{(i)\top}_{t}]\|\asymp\breve{\rho}_{n}$ with some sparsity parameter $\breve{\rho}$. If we further suppose there exists a factor $\breve{\tilde{\rho}}>0$ to describe the magnitude of the change of $\{X^{(i)}_{t}\}_{t\in[T]}$ such that $\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbb{E}[(\mathbf{O}X^{(i)}_{t_{1}}-X^{(i)}_{t_{2}})(\mathbf{O}X^{(i)}_{t_{1}}-X^{(i)}_{t_{2}})^{\top}]\|\asymp\breve{\rho}_{n}\breve{\tilde{\rho}}$ for $t_{1}\neq t_{2}$, the similar analysis as Proposition 1 and Proposition 2 can be derived.
@@ 5.1 Error in mirror estimate
We demonstrate the theoretical results for the estimated mirror in Theorem 4 with numerical simulation. Theorem 4 states a result for any fixed dynamic network $i\in[m]$. For ease of exposition, in this section we will fix a value of $i\in[m]$ and thereby drop the index $i$ from our matrices. We generate the true latent positions $\{\mathbf{X}_{t}\}$ with the following random walk process. Recall that the true latent position of vertex $s\in[n]$ in time point $t\in[T]$ is the $s$th row of $\mathbf{X}_{t}$. Let $d=1$ and denote the track of the latent position of vertex $s\in[n]$ as $\{X^{<s>}_{t}\in\mathbb{R}\}_{t\in[T]}$. Let $\tilde{c}\geq 0$ and $\delta_{T}>0$ be two constants satisfying $\tilde{c}+\delta_{T}T\leq 1$. For a fixed $p\in(0,1)$ and for all $s\in[n]$, we generate $\{X^{<s>}_{t}\in\mathbb{R}\}_{t\in[T]}$ independently as follows: $X^{<s>}_{0}=\tilde{c}$, and for $t\in[T]$,
$X^{<s>}_{t}=\begin{cases}X^{<s>}_{t-1}+\delta_{T}&\text{with probability }p,\\ X^{<s>}_{t-1}&\text{with probability }1-p.\end{cases}$ (5.1)
When we set $\delta_{T}=(1-\tilde{c})/T$, an observation is that, the dynamic network with the latent positions generated from this model has $\mathbf{B}$ with $\lambda_{1}(\mathbf{B})\asymp T$ and $\lambda_{k}(\mathbf{B})\lesssim 1$ for $k=2,\dots,T$; see Chen et al. (2024) for more details, and then we know $r=1$. We demonstrate the error rate of $\hat{\mathbf{M}}$ in Theorem 4 under this setting. More specifically, we fix $\tilde{c}=0.1,p=0.4$, and either fix $n=100$ and vary $T\in\{20,30,40,50,70,100\}$ or fix $T=10$ and vary $n\in\{50,100,200,400,800\}$. The estimation error of $\hat{\mathbf{M}}$ is evaluated using $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}\mathbf{O}-\mathbf{M}\|_{F}$, and the results are summarized in Figure 3. For this setting, according to Theorem 4 and Proposition 1, if we only vary $T$, we have $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}\mathbf{O}-\mathbf{M}\|_{F}\lesssim 1$, and if we only vary $n$, we have $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}\mathbf{O}-\mathbf{M}\|_{F}\lesssim n^{-1/2}$. We see the changes of $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}\mathbf{O}-\mathbf{M}\|_{F}$ in Figure 3 comply with the theoretical error rates obtained from Theorem 4 and Proposition 1.
@@ 5.2 Clustering performance
We perform a numerical experiment to show the clustering performance of DNCMD. We consider $K=2$ clusters of the dynamic networks generated with Eq. (5.1). These two clusters each have $20$ dynamic networks, and thus we need to cluster $m=2\times 20=40$ dynamic networks. For the dynamic networks, we fix the number of time points $T=50$, and the middle time point $t=25$ is a change point of the probability of random walk $p$. More specifically, for cluster $1$, we set $p=0.45$ for $t=1,\dots,25$ and set $p=0.55$ for $t=26,\dots,50$. And for cluster $2$, we set $p=0.55$ for $t=1,\dots,25$ and set $p=0.45$ for $t=26,\dots,50$. According to Theorem 5, Proposition 2, and Theorem 7, we know DNCMD achieves exact recovery, and it means that for $n$ large enough DNCMD can recover the true clusters with high probability. We vary $n\in\{20,30,40,80,120,200\}$ to see the clustering performance of DNCMD.
To measure the clustering accuracy, we compute the adjusted rank index (ARI) for the similarity between the estimated cluster labels and the true labels. The higher the ARI value, the closer the empirical clustering result and the true label assignment are to each other. ARI ranges from 1-1 to $1$, where $1$ indicates perfect agreement between the empirical clustering result and the true label assignment, and 0 indicates a random agreement. Since there is no existing clustering algorithm specifically designed for such dynamic network clustering problems, we compare the results of DNCMD with those of the classical clustering algorithm, k-means. More specifically, we consider using k-means to directly cluster the original adjacency matrices $\{\mathbf{A}^{(i)}\}$, or using k-means to cluster the estimated low-rank latent position matrices $\{\hat{\mathbf{X}}^{(i)}\}$. We run $100$ independent Monte Carlo replicates and summarize the result in Figure 4. Figure 4 shows that, when $n$ is large enough (in this example, $n$ only needs to exceed about $100$), DNCMD always has the ARI very close to $1$ and can recover the true clusters perfectly. And the k-means algorithm applied to the original adjacency matrices or the latent position matrices is not very effective and stable for this problem.
@@ 6 Real Data Experiments
In this section, we demonstrate the use of DNCMD on Drosophila larval connectome data to analyze edge functions in Section 6.1, and cluster the trade dynamic networks to analyze the similarities and differences in trade evolution across different food products in Section 6.2.
6.1 Drosophila larval connectome data
The complete synaptic-resolution connectome of the Drosophila larval brain was recently completed Winding et al. (2023), enabling the generation of biologically realistic models of its neural circuit based on known anatomical connectivity Eschbach et al. (2020). Training such realistic models in simulations provides significant computational power and flexibility, enabling us to study and enhance our understanding of how real animal brains operate. This approach has been applied in recent papers Acharyya et al. (2024) to investigate the operational mechanisms of the Drosophila larval brain.
The connectivity-constrained model proposed in Eschbach et al. (2020); Jiang and Litwin-Kumar (2021) can be used to generate a recurrent network of the larval mushroom body with feedback neurons, where the connectome can be constrained, and in this process, a series of stimuli can be delivered to the network; see Eschbach et al. (2020); Jiang and Litwin-Kumar (2021) for more details. Therefore, by analyzing the dynamic networks generated by this model after the removal of particular edges between neurons, we can indirectly study the function of these edges.
In this experiment, we consider the process of learning a new association between a conditioned stimulus (CS) (for example, an odor) and rewards or punishments (unconditioned stimuli, US), which serve as reinforcement. We simulate the activity of the dynamic neuron network for a total of $T=160$ time points, and at $t=16$, a random odor is delivered to the neurons in the mushroom body (CS1), followed by a reward delivered at $t=20$. We set all stimuli to last for $3$ time points. After this initial CS of odor with the reward, the odor is also delivered again without the reward at $t=80$ (CS2) and $t=140$ (CS3). The association may be weakened by exposure to the same CS without reinforcement during CS2, so we may observe a decrease in network attraction from CS2 to CS3. Here, we define network attraction as the ratio of the strength of activity in neurons responsible for attraction to the strength of activity in neurons responsible for aversion. Thus, this process includes the learning of the association and its extinction. A schematic of this process can be seen in Figure 5.
Figure 5: Schematic of the process including the learning of an association and its extinction.
We applied the DNCMD algorithm to obtain the dendrogram shown in Figure 6, with the embedding dimension for vertex latent positions set to $d=5$ and the dimension for mirror set to $r=3$. In Figure 6, we can clearly see that most of the dynamic networks resulting from the removal of edge $6$ form a distinct cluster, indicating that edge $6$ plays a critical and unique role in this process.
Figure 6: Dendrogram of dynamic networks of neurons for the removal of each single edge, obtained using the DNCMD algorithm. The labels represent the index of the removed edge (from 11 to $13$) and are shown in different colors, as indicated in the legend on the right, and for each removed edge we have $11$ replicates.
In addition, we also develop numerical measures to quantify the concentration of replicates for each label and to assess functional similarity between edges in Section A.1 of the supplementary material. These quantitative analyses further confirm that edge $6$ is special and reveal that certain edge pairs (e.g., edges $1$ and $7$, edges $5$ and $9$) exhibit similar functions.